In the laboratory, a general chemistry student measured the pH of a 0.397 M aqueous solution of methylamine, CH3NH2 to be 12.126.

Use the information she obtained to determine the Kb for this base.

Kb(experiment) =
1
Expert's answer
2019-03-25T07:30:18-0400

pH = 12.126

pOH = 14 - 12.126 = 1.874

[OH¯] = 10¯pOH = 10¯1.874 = 1.34 x 10-2 M

Kb = ([CH3NH3+]*[OH-])/[CH3NH2] = (0.00134*0.00134)/(0.397) = 4.5 x 10-4

Answer: Kb = 4.5 * 10-4